A six-sided die is biased in such a way that even numbers are three times as likely as odd numbers.what is the probability of observing at least a four on one roll of this die?
1
Expert's answer
2021-09-03T07:51:34-0400

Let "P(X=1)=P(X=3)=P(x=5)=a."

Then "P(X=2)=P(X=4)=P(x=6)=3a."



"a+3a+a+3a+a+3a=1"

"a=\dfrac{1}{12}"

Hence

"P(X=4)=3(\dfrac{1}{12})=\dfrac{1}{4}"